How to draw a rectangle with an area of 5 square units and a perimeter of 21 units?
Answer by TimothyLamb(4379)

xy = 5
2x + 2y = 21
2x = 21 - 2y
x = 21/2 - y
---
xy = 5
(10.5 - y)y = 5
(10.5)y - yy - 5 = 0
-yy + 10.5y - 5 = 0
---
the above quadratic equation is in standard form, with a=-1, b=10.5, and c=-5
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
-1 10.5 -5
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two x-intercepts), of the quadratic are:
x = 0.5
x = 10
---
answer:
length = 10 units
width = 1/2 unit
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php